Skip to main content

Service status

CURRENT ISSUES

On-going issues

  • Lynda.com continues to be unavailable for most users, this started on Wednesday 20 February and is on-going. We are working with LinkedIn Learning Technical Support to migrate users learning data from Lynda.com and make the LinkedIn Learning Service available for our users. Apologies for the inconvenience this extended outage is causing. 

 

PLANNED MAINTENANCE - Weekly every TUESDAY Morning

Maintenance window 06.00 am to 10.00 am

  • Tuesday 26 March 2019  Novicom (drives the information screens around the College) server upgrade - no user impact expected
  • Tuesday 26 March 2019  Test network failover from BMA to Gordon Square/Malet Street

 

PLANNED MAINTENANCE - future

13th and 14th April - Egmont house - NO network access

  • Network changes are required to be made at Egmont House affecting all network connected devices.
  • Staff will not be able to used PC's, Laptops, WiFi or printing over the weekend.
  • Staff will not be able to remotely connect to their PC's either.

 

Monday 01 April 2019 09.00am until Tuesday 02 April 2019 12.00pm - Upgrades to Event Management Systems including Downtime

  • From 09.00am on Monday 01 April 2019 until around 12.00pm the following day it will not be possible to add or otherwise update events using the Events Management System in My Birkbeck for Staff.
  • For approximately 1 hour during this window, likely on the morning of Tuesday 02 April 2019 events will not be visible on the web. Every effort will be made to keep this downtime to a minumum
  • Following this upgrade work, if it is successful, significant improvements to the Events Management System will be available. Details of these improvements will be circulated via email, and in My Birkbeck for Staff.
  • Lastly, this upgrade work will include a replacement system for managing and booking on Study Skills Workshops.

 

RESOLVED and COMPLETED

RESOLVED - 09.00AM FRIDAY 22 MARCH 07.00AM MAJOR NETWORK INCIDENT

  • An engineer from UCL has assisted, and all network services are now up and running.

resolved - Thursday 21 march - virtualisation platform hardware failure

  • Hardware issue arising Thursday 14 March on the virtualisation platform - hardware has been replaced by an HP engineer and this is now resolved.

REsolved - WEDNESDAY 20 MARCH 2019 07.50AM - VDI SERVICES AFFECTING LIBRARY AND USS ARE UNAVAILABLE

  • Initial investigations indicate a software corruption.
  • This is now resolved. 

RESOLVED WEDNESDAY 20 MARCH 2019 08.50am - MAIN WEB SITE IS UNAVAILABLE

  • Resolved at 09.25am.
  • Default "home" pages, including main home page not performing as expected.
  • Departmental areas of the site remained operational.

COMPLETED - Wednesday 06 MArch 22 FEBRUARY 2019 - PLANNED MAINTENANCE AFFECTING ROOM BOOKINGS, STUDY SKILLS WORKSHOPS AND SOME OTHER FORMS

  • Planned maintenance on the web server that hosts the following systems:
    • Room bookings
    • Study Skills Workshop bookings
    • Design requests
    • Alumni contact detail updates
    • Contact us form for the Department of Organizational Psychology
    • Requests for items from Library stores
  • This work is now completed. 

COMPLETED - HORIZON VIEW VDI - FRIDAY 22 FEBRUARY 2019

This work is now completed. SSL certificates were replaced on connection servers to maintain secure communications with the platform.

COMPLETED - FRIDAY 22 FEBRUARY 2019 - TELEPHONE PROBLEM IN EGMONT HOUSE

  • There was a problem with the telephone switch in Egmont House between midnight and 10.30am on Friday 22 February 2019. The impact was short dialing (4 digits) to some extensions was unavailable, the service is now fully operational.

COMPLETED - TUESDAY 12 FEBRUARY 2019 - MOODLE UNAVAILABLE FOR MAINTENANCE

Moodle was unavailable between 07.00pm and 09.30pm on Tuesday 19 February 2019. We apologise for the inconvenience caused and are waiting for a report from our hosting service for the reason. 

COMPLETED - TUESDAY 12 FEBRUARY 2019 - Exchange maintenance work

Maintenance work on our Exchange Email service was successfully completed. 

This work was to alleviate performance problems that have been reported by Outlook users on the Eduroam network.

RESOLVED - FRIDAY 25 JANUARY 2019 - problems logging in to Virtual Desktop issues

  • Virtual Desktop (VDI) services have now been restored and are fully operational as at 11:50am on 25/1/2019.

COMPLETE - THURSDAY 6 DECEMBER 2018 - MANDATED STAFF PASSWORD CHANGE

  • All Birkbeck Staff are required to change their password by 12.00 midday Friday 7 December 2018.
  • This also applies to other users – eg visiting academics, contractors, partners etc – but NOT students, applicants, prospective students and alumni.
  • Following midday Friday 7 December, if you have not changed your password since 09.00am on Monday 3 December 2018 your password will be forcibly changed to a random value. Where this occurs you will be required to complete password recovery to regain access your account.
  • You can change your password using the self-service password facility.
  • Following midday 7 December 2018 we will scramble your account password if it remains unchanged. If you have a personal email address recorded in our systems you will be able to perform a self-service password reset. Otherwise, you will be required to contact IT Services who will verify your identity and recover your account.

resolved - 26 Nov 2018 09.00am - SPAM MAIL - ALLEGEDLY FROM S.W*****N@BBK.AC.UK

  • An email allegedly from s.w*****n@bbk.ac.uk with the Subject: Update has been circulated.

  • The email has a link called "Pictures" - it is malicious. Please delete the email.

  • If you have accidentally clicked the Pictures link, please contact IT Services on extension 6543 or 020 7631 6543.

resolved - 23 Nov 2018 11.00am - INCOMING EMAIL FAULT

  • (23 November 2018 10.00am): Please be aware that as of ~07.00am this morning there is a problem with mail coming in to the College from outside.

  • Specifically the problem is with 'smtp.bbk.ac.uk'.

  • This has been resolved - this appears to be a denial of service attack.